# OPERA: A Reinforcement Learning–Enhanced Orchestrated Planner-Executor Architecture for Reasoning-Oriented Multi-Hop Retrieval

### v2

@@ 1 Introduction
The ability to solve complex problems is a core aspect of intelligence, and within Retrieval-Augmented Generation (RAG), reasoning-centric retrieval provides an effective means to address these tasks in the post-training era. The concurrent improvement of Large Language Models (LLMs) (brown2020language; devlin2019bert) and dense retrieval systems (karpukhin2020dense; khattab2020colbert) has propelled the evolution of RAG. The traditional RAG follows a retrieve-then-reason paradigm (lee2019latent; lewis2020retrieval), which has now been widely optimized into a multi-stage pipeline including query rewriting, document retrieval, document filtering, and answer generation (izacard2020leveraging). However, despite significant progress, effectively orchestrating these capabilities remains challenging, as existing approaches struggle with the demands of multi-hop reasoning. Consider a query such as, “What was the previous occupation of the person who succeeded the founder of the company that acquired WhatsApp?” Such questions demand not merely retrieving documents, but orchestrating a precise sequence of retrieval and reasoning steps where each operation depends critically on its predecessors.
Current approaches face several key challenges. First, existing solutions have limited reasoning-oriented planning. While planner-first models like PlanRAG (lee2024planrag) and REAPER (reaper2024) introduce upfront planning mechanisms, their static plans cannot dynamically adapt to unforeseen challenges during retrieval. Second, most methods lack effective reasoning-aware retrieval. Even adaptive methods like Adaptive-RAG (jeong2024adaptive) and AT-RAG (zhang2024atrag), which adjust retrieval strategy based on query complexity, lack the fine-grained, reasoning-driven query reformulation needed for complex multi-hop scenarios. Third, current systems provide inadequate reasoning-guided filtering. Even when retrieval fetches golden documents, they are often buried within noisy top-K results. While iterative approaches like ReAct (yao2022react) attempt to address this through reasoning-action loops, and recent methods such as BGE (ke2024bridging) show promise in bridging retriever-LLM preferences, their effectiveness remains limited by indirect reward signals that fail to capture the nuanced reasoning required for effective filtering. These limitations persist because advanced enhancement strategies—spanning SFT, preference optimization, and RL—are insufficient or misaligned, often leading to goal misalignment between modules (qi2023fine; wang2023instruction). These issues stem from fundamental weakness in the coupling between retrieval and reasoning, preventing full utilization of modern LLMs and dense retrievers’ capabilities.
@@ 2 Related Work
RAG. To mitigate hallucination, Retrieval-Augmented Generation (RAG) was introduced to ground outputs in external knowledge (lewis2020retrieval). The initial “retrieve-then-read” paradigm (lee2019latent), however, proved insufficient for multi-hop reasoning due to its static pipeline (trivedi2022musique; arabzadeh2021predicting; luan2021sparse). The field has since evolved toward more dynamic retrieval, for instance by routing queries based on complexity (jeong2024adaptive) or introducing explicit planning (lee2024planrag). While these methods improve upon static RAG, they primarily optimize the retrieval act itself, rather than the overarching reasoning strategy. OPERA is distinct in its hierarchical approach: a specialized Goal Planning Module (GPM) governs high-level strategy, while an agentic Reason-Execute Module (REM) handles tactical execution, including fine-grained analysis and adaptive query reformulation.
@@
Chain-of-Thought. Chain-of-Thought (CoT) methods (wei2022chain; yao2023tree) and agentic frameworks like ReAct (yao2022react) and IRCoT (trivedi2023interleaving) established the importance of decomposing problems and interleaving reasoning with actions. However, these frameworks rely on a single, general LLM for high-level planning and low-level execution, which can compromise reliability. While multi-agent systems like MetaGPT (hong2024metagpt) assign distinct roles, they typically use generalist models. OPERA advances by employing an asymmetric architecture, pairing the strategic Goal Planning Module (GPM) with the agentic Reason-Execute Module (REM) to separate strategic and tactical concerns.
@@
Reinforcement Learning. A key aspect of our work is the training of our specialized planner and executor. Many frameworks utilize on-policy algorithms like Proximal Policy Optimization (PPO); however, PPO often struggles with the large action spaces and sparse rewards common in RAG (stiennon2020learning; fernandezoLooria2022survey; ramamurthy2023reinforcement), making training unstable and sample-inefficient. To address this, preference-based optimization has gained traction. While Direct Preference Optimization (DPO) (rafailov2023direct) has become a standard for learning from binary preferences (chosen >> rejected), it is ill-suited for more nuanced reward signals (ivison2024unpacking). Our training methodology generates fine-grained scalar scores reflecting the quality of a plan or an execution step. Using DPO would discard this rich information by compressing it into a binary signal. To fully leverage these scalar rewards within our multi-agent setting, we build upon Group Relative Policy Optimization (GRPO) (shao2024grpo) and introduce our own variant: Multi-Agents Progressive Group Relative Policy Optimization (MAPGRPO). Unlike standard GRPO, MAPGRPO is specifically designed for our staged training protocol, enabling fine-grained credit assignment (papangelis2019collaborative) and ensuring coordinated optimization across the distinct roles of the GPM and REM agents.
@@ 3 Method
We propose MAPGRPO, a novel variant of Group Relative Policy Optimization (GRPO) (shao2024grpo).
@@ 4 Experimental
Datasets. We evaluate OPERA on three multi-hop reasoning benchmarks: HotpotQA (yang2018hotpotqa) (90K questions), 2WikiMultiHopQA (ho2020constructing) (150K questions), and Musique (trivedi2022musique) (25K questions). For out-of-domain evaluation, we use NQ (kwiatkowski2019natural) and MultiHopRAG (tang2024multihop).
Implementation Details. We use Qwen2.5-7B-Instruct***https://huggingface.co/Qwen/Qwen2.5-7B-Instruct for Plan and Analysis-Answer agents and other baseline’s backbone, Qwen2.5-3B-Instructhttps://huggingface.co/Qwen/Qwen2.5-3B-Instruct for the Rewrite agent, and BGE-M3 (chen2024bge) as our dense retriever with top-5 document retrieval. For pre-scored dataset construction, we utilize the DeepSeek R1 APIDeepSeek R1 is accessed via the official API service provided by DeepSeek (https://platform.deepseek.com). as the data generation model. Training details are in Appendix A.2.
Baselines. We compare OPERA against methods in four categories: Naive: (1) Qwen2.5-7B (No Retrieval); (2) Single-Step RAG. CoT: (3) IRCoT (trivedi2023interleaving); (4) OPERA (CoT Only). SFT: (5) Adaptive-RAG (jeong2024adaptive). RL: (6) BGM (ke2024bridging); (7) OPERA (MAPGRPO). All SFT and RL methods are trained on mixed datasets, with Plan Agent utilizing pre-scored dataset $\mathcal{D}_{\text{scored}}$ for high-score sample selection (Appendix A.6). Baselines use defaults with Faiss (douze2024faiss).
Evaluation Metrics. We report EM (%) (exact match), F1 (%) (token-level overlap), Steps (average reasoning steps), Latency (processing time), and Success Rate (execution completion rate).
@@
Method HotpotQA 2WikiMultiHopQA Musique Type
EM (%) F1 (%) EM (%) F1 (%) EM (%) F1 (%)
Qwen2.5-7B (No Retrieval) 18.5 26.8 16.2 23.7 4.1 9.1 Single LLM
Single-Step RAG 31.5 44.2 25.9 37.6 14.1 18.4 Naive RAG
IRCoT (trivedi2023interleaving) 42.7 54.8 43.3 56.2 18.8 23.9 CoT
OPERA (CoT) 44.9 58.5 42.3 50.7 21.2 26.4 CoT
Adaptive-RAG (jeong2024adaptive) 45.7 56.9 30.1 39.3 24.3 35.7 SFT
BGM (ke2024bridging) 41.5 53.8 44.3 55.8 19.6 26.8 RL
Table 1 shows OPERA’s performance across all benchmarks.
@@
Figure 3: OPERA’s runtime dynamics. (Left) Agent call intensity and question completion rate over processing steps. (Right) Attention visualization across query token types and processing stages.
Figure 4: Training dynamics for MAPGRPO. (Top) Average reward curves show stable learning. (Bottom) Plots confirm controlled policy divergence (left) and significant gradient variance reduction from expert injection (right).
Figure 5: Reward distribution evolution during MAPGRPO training for Plan and Analysis-Answer agents.
@@
Trajectory and Training Visualization
Runtime Dynamics and Attention Flow. Figure 3 shows OPERA’s decision-making and execution patterns. The left panel shows agent activation over time: the Plan Agent (blue) initiates question cycles, the Analysis-Answer Agent (green) performs reasoning, and the Rewrite Agent (orange) activates upon retrieval failures. The black trajectory tracks the cumulative question completion rate, demonstrating performance across three questions of varying complexity. The right panel illustrates attention evolution across processing stages—from entity-focused planning to relation-aware analysis and context-integrated refinement.
Stable Convergence and Reward Evolution. MAPGRPO enables stable and efficient training with consistent performance improvements. Training dynamics are shown in Figures 4 and 5. Figure 4 shows typical RL dynamics: initial instability (epochs 1–5) followed by gradual improvement with occasional dips, most evident in the conditionally activated Rewrite Agent. The shaded region in the bottom-right panel highlights the Expert Injection Impact, showing significant variance reduction for the 7B model with expert injection compared to its counterpart without, effectively improving training stability and confirming our theoretical analysis.
Figure 5 illustrates this learning progress by visualizing the evolution of reward distributions. For both the Plan and Analysis-Answer agents, the distributions progressively shift toward higher reward values and become more concentrated as training advances. This suggests that the agents improve consistently rather than just on average.
@@
Out-of-Domain Evaluation. Table 3 shows how training methods affect generalization. On NQ—single-hop QA over full Wikipedia pages where planning is unnecessary—MAPGRPO achieves 36.6% EM while SFT drops to 19.5% (from 23.1% CoT baseline). MAPGRPO preserves OPERA’s flexibility, allowing it to bypass planning for single-hop queries and use the Analysis-Answer Agent’s training on (sub-question, document, answer) tuples to handle long documents. SFT appears to overfit to multi-hop patterns, attempting decomposition even when not needed. On MHRAG, which has multi-hop structure similar to our training data, all methods improve, with MAPGRPO reaching 55.7% EM. RL methods perform well on both single and multi-hop tasks while SFT performs worse on single-hop tasks. This suggests that trajectory-based optimization with diverse rewards helps OPERA adapt its reasoning strategy to different tasks, while supervised learning tends toward consistent multi-hop behavior.
@@ Appendix A Appendix
aQwen2.5 series (yang2024qwen2.5). bLLama3.1 series (grattafiori2024llama3). cQwen3 (yang2025qwen3).
@@
Candidate Generation Process. For each question $q$ with ground-truth answer $a^{*}$ and supporting facts $facts^{*}$, we generate candidates using the DeepSeek R1 API service§§§We use DeepSeek R1 via the official commercial API for data generation.:
@@
Planning Inefficiencies. Many questions allow multiple valid decomposition paths, but OPERA lacks explicit optimization for path efficiency and may select suboptimal reasoning chains. Several research directions could address these limitations. Developing adaptive decomposition methods with path-efficiency optimization would improve planning effectiveness. Creating specialized datasets for long-chain reasoning would enhance performance on complex multi-hop questions. Hybrid architectures that maintain modularity while reducing computational overhead could make the system more practical for real-time use. Advanced retrieval techniques like Beam Retrieval (zhang2024endtoend) could explore multiple retrieval paths simultaneously, improving robustness for ambiguous queries.
@@
The prompt templates for each agent are presented in Figure 9, 10 and 11.

### v3

@@ 1 Introduction
The ability to solve complex problems is a core aspect of intelligence, and within Retrieval-Augmented Generation (RAG), reasoning-centric retrieval provides an effective means to address these tasks in the post-training era. The concurrent improvement of Large Language Models (LLMs) (Brown et al. 2020; Devlin et al. 2019) and dense retrieval systems (Karpukhin et al. 2020; Khattab and Zaharia 2020) has propelled the evolution of RAG. The traditional RAG follows a retrieve-then-reason paradigm (Lee et al. 2019; Lewis et al. 2020), which has now been widely optimized into a multi-stage pipeline including query rewriting, document retrieval, document filtering, and answer generation (Izacard and Grave 2021). However, despite significant progress, effectively orchestrating these capabilities remains challenging, as existing approaches struggle with the demands of multi-hop reasoning. Consider a query such as, “What was the previous occupation of the person who succeeded the founder of the company that acquired WhatsApp?” Such questions demand not merely retrieving documents, but orchestrating a precise sequence of retrieval and reasoning steps where each operation depends critically on its predecessors.
Current approaches face several key challenges. First, existing solutions have limited reasoning-oriented planning. While planner-first models like PlanRAG (Lee et al. 2024) and REAPER (Joshi et al. 2024) introduce upfront planning mechanisms, their static plans cannot dynamically adapt to unforeseen challenges during retrieval. Second, most methods lack effective reasoning-aware retrieval. Even adaptive methods like Adaptive-RAG (Jeong et al. 2024) and AT-RAG (Rezaei et al. 2024), which adjust retrieval strategy based on query complexity, lack the fine-grained, reasoning-driven query reformulation needed for complex multi-hop scenarios. Third, current systems provide inadequate reasoning-guided filtering. Even when retrieval fetches golden documents, they are often buried within noisy top-K results. While iterative approaches like ReAct (Yao et al. 2023b) attempt to address this through reasoning-action loops, and recent methods such as BGE (Ke et al. 2024) show promise in bridging retriever-LLM preferences, their effectiveness remains limited by indirect reward signals that fail to capture the nuanced reasoning required for effective filtering. These limitations persist because advanced enhancement strategies—spanning SFT, preference optimization, and RL—are insufficient or misaligned, often leading to goal misalignment between modules (Qi et al. 2024; Wang et al. 2023). These issues stem from fundamental weakness in the coupling between retrieval and reasoning, preventing full utilization of modern LLMs and dense retrievers’ capabilities.
@@ 2 Related Work
RAG. To mitigate hallucination, Retrieval-Augmented Generation (RAG) was introduced to ground outputs in external knowledge (Lewis et al. 2020). The initial “retrieve-then-read” paradigm (Lee et al. 2019), however, proved insufficient for multi-hop reasoning due to its static pipeline (Trivedi et al. 2022; Arabzadeh et al. 2021; Luan et al. 2021). The field has since evolved toward more dynamic retrieval, for instance by routing queries based on complexity (Jeong et al. 2024) or introducing explicit planning (Lee et al. 2024). While these methods improve upon static RAG, they primarily optimize the retrieval act itself, rather than the overarching reasoning strategy. OPERA is distinct in its hierarchical approach: a specialized Goal Planning Module (GPM) governs high-level strategy, while an agentic Reason-Execute Module (REM) handles tactical execution, including fine-grained analysis and adaptive query reformulation.
@@
Chain-of-Thought. Chain-of-Thought (CoT) methods (Wei et al. 2022; Yao et al. 2023a) and agentic frameworks like ReAct (Yao et al. 2023b) and IRCoT (Trivedi et al. 2023) established the importance of decomposing problems and interleaving reasoning with actions. However, these frameworks rely on a single, general LLM for high-level planning and low-level execution, which can compromise reliability. While multi-agent systems like MetaGPT (Hong et al. 2024) assign distinct roles, they typically use generalist models. OPERA advances by employing an asymmetric architecture, pairing the strategic Goal Planning Module (GPM) with the agentic Reason-Execute Module (REM) to separate strategic and tactical concerns.
@@
Reinforcement Learning. A key aspect of our work is the training of our specialized planner and executor. Many frameworks utilize on-policy algorithms like Proximal Policy Optimization (PPO); however, PPO often struggles with the large action spaces and sparse rewards common in RAG (Stiennon et al. 2020; Uc-Cetina et al. 2022; Ramamurthy et al. 2023), making training unstable and sample-inefficient. To address this, preference-based optimization has gained traction. While Direct Preference Optimization (DPO) (Rafailov et al. 2023) has become a standard for learning from binary preferences (chosen >> rejected), it is ill-suited for more nuanced reward signals (Ivison et al. 2024). Our training methodology generates fine-grained scalar scores reflecting the quality of a plan or an execution step. Using DPO would discard this rich information by compressing it into a binary signal. To fully leverage these scalar rewards within our multi-agent setting, we build upon Group Relative Policy Optimization (GRPO) (Shao et al. 2024) and introduce our own variant: Multi-Agents Progressive Group Relative Policy Optimization (MAPGRPO). Unlike standard GRPO, MAPGRPO is specifically designed for our staged training protocol, enabling fine-grained credit assignment (Papangelis et al. 2019) and ensuring coordinated optimization across the distinct roles of the GPM and REM agents.
@@ 3 Method
We propose MAPGRPO, a novel variant of Group Relative Policy Optimization (GRPO) (Shao et al. 2024).
@@
Method HotpotQA 2WikiMultiHopQA Musique Type
EM (%) F1 (%) EM (%) F1 (%) EM (%) F1 (%)
Qwen2.5-7B (No Retrieval) 18.5 26.8 16.2 23.7 4.1 9.1 Single LLM
Single-Step RAG 31.5 44.2 25.9 37.6 14.1 18.4 Naive RAG
IRCoT (Trivedi et al. 2023) 42.7 54.8 43.3 56.2 18.8 23.9 CoT
OPERA (CoT) 44.9 58.5 42.3 50.7 21.2 26.4 CoT
Adaptive-RAG (Jeong et al. 2024) 45.7 56.9 30.1 39.3 24.3 35.7 SFT
BGM (Ke et al. 2024) 41.5 53.8 44.3 55.8 19.6 26.8 RL
@@ 4 Experimental
More training and experiment details are in Appendix A.2. Datasets. We evaluate OPERA on three multi-hop reasoning benchmarks: HotpotQA (Yang et al. 2018) (90K questions), 2WikiMultiHopQA (Ho et al. 2020) (150K questions), and Musique (Trivedi et al. 2022) (25K questions). For out-of-domain evaluation, we use NQ (Kwiatkowski et al. 2019) and MultiHopRAG (Tang and Yang 2024).
Implementation Details. We use Qwen2.5-7B-Instruct (Yang et al. 2024) for Plan and Analysis-Answer agents and other baseline’s backbone, Qwen2.5-3B-Instruct (Yang et al. 2024) for the Rewrite agent, and BGE-M3 (Chen et al. 2024) as our dense retriever with top-5 document retrieval. For pre-scored dataset construction, we utilize the DeepSeek R1 (DeepSeek-AI et al. 2025) API as the data generation model.
@@
Baselines. We compare OPERA against methods in four categories: Naive: (1) Qwen2.5-7B (No Retrieval); (2) Single-Step RAG. CoT: (3) IRCoT (Trivedi et al. 2023); (4) OPERA (CoT Only). SFT: (5) Adaptive-RAG (Jeong et al. 2024). RL: (6) BGM (Ke et al. 2024); (7) OPERA (MAPGRPO). All SFT and RL methods are trained on mixed datasets, with Plan Agent utilizing pre-scored dataset for high-score sample selection. Baselines use defaults with Faiss (Douze et al. 2024).
Evaluation Metrics. We report EM (%) (exact match), F1 (%) (token-level overlap), Steps (average reasoning steps), Latency (processing time), and Success Rate (execution completion rate). Table 1 shows OPERA’s performance across all benchmarks.
@@
Figure 3: OPERA’s runtime dynamics. (Top) Agent call intensity and question completion rate over processing steps. (Bottom) Attention visualization across query token types and processing stages.
Figure 4: Training dynamics for MAPGRPO across three training stages. Top row shows average reward curves for (top-left) Plan Agent with expert samples and (top-right) Analysis-Answer Agent. Bottom row presents (bottom-left) Rewrite Agent training dynamics and (bottom-right) policy gradient variance reduction, with the shaded region highlighting the expert injection impact.
@@
Trajectory and Training Analysis
Runtime Dynamics and Attention Flow. Figure 3 shows OPERA’s decision-making and execution patterns. The top panel shows agent activation over time: the Plan Agent (blue) initiates question cycles, the Analysis-Answer Agent (green) performs reasoning, and the Rewrite Agent (orange) activates upon retrieval failures. The black trajectory tracks the cumulative question completion rate, demonstrating performance across three questions of varying complexity. The bottom panel illustrates attention evolution across processing stages—from entity-focused planning to relation-aware analysis and context-integrated refinement.
Stable Convergence and Reward Evolution. In reward-driven GRPO, each training step samples eight diverse gold/silver candidates per input, enabling contrastive learning that steers the policy toward better outputs. These signals help agents reach or even surpass $C_{\text{best}}$ (Eq. 11). MAPGRPO enables stable and efficient training with consistent performance improvements. Figure 4 shows typical RL characteristics: initial instability in the early steps, followed by progressive improvement with occasional dips, particularly visible in the Rewrite Agent due to its conditional activation. Our expert demonstration strategy (Expert Injection Impact) is highly effective at reducing policy gradient variance, a key factor in training stability. The shaded region in the bottom-right panel highlights this effect, showing a significant variance reduction for the 7B model with expert injection compared to the no-expert variant, confirming our theoretical analysis. The Rewrite Agent is more unstable because it activates only on retrieval failures, yielding sparse rewards; retriever limitations further make Golden-Documents inherently difficult to obtain. Theory (Eq. 6) guarantees only local (not global) convergence and thus no performance guarantee (e.g., Musique EM remains <<40% in Table 1).
@@
Out-of-Domain Evaluation. Table 3 shows how training methods affect generalization. On NQ—single-hop QA over Wikipedia pages where planning is unnecessary—MAPGRPO achieves 36.6% EM while SFT drops to 19.5% (from 23.1% CoT baseline). MAPGRPO preserves OPERA’s flexibility, allowing it to bypass planning for single-hop queries and use the Analysis-Answer Agent’s training on (sub-question, document, answer) tuples to handle long documents. SFT overfits to multi-hop patterns, attempting decomposition even when not needed. On MHRAG, which has multi-hop structure similar to training data, all methods improve, with MAPGRPO reaching 55.7% EM. RL methods perform well on both single- and multi-hop tasks, whereas SFT performs worse on single-hop tasks. This suggests that trajectory-based optimization enables adaptive reasoning, while SFT induces rigid behavior.
@@ Appendix A Appendix
aQwen2.5 series (Yang et al. 2024). bLLama3.1 series (Grattafiori et al. 2024). cQwen3 (Yang et al. 2025).
@@
Corpus Settings. We construct a large-scale unified retrieval corpus by integrating three benchmark datasets: HotpotQA (Yang et al. 2018), 2WikiMultiHopQA (Ho et al. 2020), and MusiQue (Trivedi et al. 2022). Rather than indexing entire documents, we adopt a sentence-level indexing strategy for precise retrieval. We extract all sentence-level chunks from the development, test, and training splits of these datasets, then perform exact deduplication based on $(\text{title},\text{content})$ tuples to remove redundant information. The resulting corpus is indexed using BGE-M3 embeddings (Chen et al. 2024) with FAISS (Johnson et al. 2019) for efficient dense retrieval. Table 9 presents the corpus statistics. After deduplication, the final corpus contains 1.78 million unique document chunks, providing comprehensive coverage for multi-hop reasoning evaluation. During evaluation on a specific dataset, the retrieval system accesses the entire unified corpus, creating a realistic open-domain scenario where relevant documents must be identified among diverse distractors from all datasets.
Dataset Unique Chunks
HotpotQA-Dev 269,558
2Wiki-Dev 206,743
2Wiki-Test 205,465
2Wiki-Train 1,346,267
MusiQue-Dev 9,732
MusiQue-Train 47,380
Sum Before Cross-Dataset Dedup 2.09M (2,085,145)
Final After Cross-Dataset Dedup 1.78M (1,780,294)
Table 9: Statistics of the unified retrieval corpus showing cross-dataset deduplication effects.
@@
Candidate Generation Process. For each question $q$ with ground-truth answer $a^{*}$ and supporting facts $facts^{*}$, we generate candidates using the DeepSeek R1 API service***We use DeepSeek R1 via the official commercial API for data generation.:
@@
Planning Inefficiencies. Many questions allow multiple valid decomposition paths, but OPERA lacks explicit optimization for path efficiency and may select suboptimal reasoning chains. Several research directions could address these limitations. Developing adaptive decomposition methods with path-efficiency optimization would improve planning effectiveness. Creating specialized datasets for long-chain reasoning would enhance performance on complex multi-hop questions. Hybrid architectures that maintain modularity while reducing computational overhead could make the system more practical for real-time use. Advanced retrieval techniques like Beam Retrieval (Zhang et al. 2024) could explore multiple retrieval paths simultaneously, improving robustness for ambiguous queries.
Corpus Constraints and Domain Scope. The primary objective of our experimental design is to validate the structural rationality of the OPERA architecture and the effectiveness of the MAPGRPO training protocol within a controlled, high-density information environment. However, a primary limitation of the current evaluation is that OPERA operates within a closed-domain retrieval setting. The knowledge base is synthesized from the official document collections of the three source datasets, which ensures high information density but may not fully represent the noise and ambiguity of the open web. This restricted scope limits the system’s exposure to massive-scale entity disambiguation challenges. Future work will focus on extending OPERA to open-domain scenarios using the full Wikipedia corpus and real-time web indexing. This extension will test the architecture’s robustness in navigating more diverse knowledge distributions and answering more generalized, cross-domain questions.
@@
The prompt templates for each agent are presented in Figure 8, 9 and 10.